OpenStudy (anonymous):

You and your friend are shopping for painting supplies. Your friend wants five bottles of paint and one set of paint brushes. You want seven bottles of paint and two sets of paint brushes. Each bottle of paint costs $3.69. If you and your friend’s total cost is $51.69, what is the cost of one set of paint brushes?

OpenStudy (anonymous):

Directions: Translate each real-world situations into a multi-step equation. Solve the equation, showing all steps to solve the equation. Then write an explanation of the answer to the problem.

OpenStudy (texaschic101):

your friend -- 5b + 1p you -- 7b + 2p p = 3.69 5b + 1p + 7b + 2p = 51.69 12b + 3p = 51.69 (now sub 3.69 in for p) 12b + 3(3.69) = 51.69 12b + 11.07 = 51.69 12b = 51.69 - 11.07 12b = 40.62 b = 40.62/12 b = 3.385 check.. 12b + 11.07 = 51.69 12(3.385) + 11.07 = 51.69 40.62 + 11.07 = 51.69 51.69 = 51.69 (correct) paint brushes cost $ 3.385....rounds to 3.39